# The Silent Thought: Modeling Internal Cognition in Full-Duplex Spoken Dialogue Models via Latent Reasoning

### 18 Mar 2026 v1

@@ 3.1 Full-duplex SDLMs
Finally, an autoregressive decoder (parameterized by $\psi$) predicts the agent’s speech tokens $Y^{\mathrm{spc}}$ based on the text tokens and the hidden states $h_{t}^{l}$ from the LLM:
$\begin{split}P_{\psi}(Y^{spc}&|X,Y^{txt},h)=\\ &\prod_{t=1}^{T}P_{\psi}(Y_{t}^{spc}|Y_{<t}^{spc},Y_{\leq t}^{txt},h^{l}_{\leq t}\rangle).\end{split}$ (2)
@@ 3.2 Evidence Lower Bound
In Latent Variable Models, maximizing the marginal likelihood $\log P_{\theta}(x)$ is typically intractable due to the high-dimensional integration required over the latent space. To address this, Variational Inference introduces a parameterized distribution $q_{\phi}(z|x)$ to approximate the inaccessible true posterior (Zhang et al., 2018). The ELBO serves as a surrogate optimization objective, derived via Jensen’s Inequality to provide a functional lower bound on the log-likelihood:
$\begin{split}\text{log}P_{\theta}(x)\geq\mathbb{E}{q_{\phi}(z|x)}&[\log P_{\theta}(x|z)]-\text{KL}[q_{\phi}(z|x)||P(z)].\end{split}$ (4)
@@ 4.2 SFT for Latent Reasoning
Consequently, we can implement efficient latent reasoning SFT with teacher forcing by utilizing both the latent reasoning embedding labels and the text token labels. Upon feeding $H_{in}$ into the LLM backbone, we obtain the last hidden state $h^{l}$. This state is then passed through the LLM head and a Softmax function to predict the text logits $\hat{y}^{txt}$. By applying the argmax operation over these logits, we obtain the predicted text tokens $\hat{Y}^{txt}$. Inspired by (5), the loss function for ELBO-based latent reasoning training comprises two components. The first component is the conditional reconstruction loss, specifically the next-token prediction loss for the response segment, defined as:
@@
Here, $\mathrm{sg}[\cdot]$ denotes the stop-gradient operator. We use the vocabulary-weighting distribution produced by the Global-aware Expert model—namely, the weights used to compute the latent-reasoning embedding labels—to supervise the vocabulary-weighting distribution predicted by the LLM during the user’s speaking phase (i.e., the listening phase). The latent reasoning timing label $G$ is similarly applied to ensure that this loss is computed exclusively during the user’s speaking phase.
@@ 5.2 Training pipeline
FLAIR w/o latent thkthk 0.33 0.49 100 4.3
FLAIR w/ latent thkthk 0.39 0.46 100 4.3

### 03 Apr 2026 v2

@@ 3.1 Full-duplex SDLMs
Finally, an autoregressive decoder (parameterized by $\psi$) predicts the agent’s speech tokens $Y^{\mathrm{spc}}$ based on the text tokens and the hidden states $h^{l}$ from the LLM:
$\begin{split}P_{\psi}(Y^{spc}&|X,Y^{txt},h^{l})=\\ &\prod_{t=1}^{T}P_{\psi}(Y_{t}^{spc}|Y_{<t}^{spc},Y_{\leq t}^{txt},h^{l}_{\leq t}\rangle).\end{split}$ (2)
@@ 3.2 Evidence Lower Bound
In Latent Variable Models, maximizing the marginal likelihood $\log P_{\theta}(x)$ is intractable due to the high-dimensional integration required over the latent space. To address this, Variational Inference introduces a parameterized distribution $q_{\phi}(z|x)$ for latent variable $z$ to approximate the true posterior (Zhang et al., 2018). The ELBO serves as a surrogate optimization objective, derived via Jensen’s Inequality to provide a functional lower bound on the log-likelihood:
$\begin{split}\text{log}P_{\theta}(x)\geq\mathbb{E}_{q_{\phi}(z|x)}&[\log P_{\theta}(x|z)]-\text{KL}[q_{\phi}(z|x)||P(z)].\end{split}$ (4)
@@ 4.2 SFT for Latent Reasoning
Consequently, we can implement efficient latent reasoning SFT with teacher forcing by utilizing both the latent reasoning embedding labels and the text token labels. Upon feeding $H_{in}$ into the LLM backbone, we obtain the last hidden state $h^{l}$. This state is then passed through the LLM head to predict the text logits $\hat{y}^{txt}$. By applying the argmax operation over these logits, we obtain the predicted text tokens $\hat{Y}^{txt}$. Inspired by (5), the loss function for ELBO-based latent reasoning training comprises two components. The first component is the conditional reconstruction loss, specifically the next-token prediction loss for the response segment, defined as:
@@
Here, $\mathrm{sg}[\cdot]$ denotes the stop-gradient operator. We use the vocabulary-weighting distribution produced by the Global-aware Expert model, namely, the weights used to compute the latent-reasoning embedding labels, to supervise the vocabulary-weighting distribution predicted by the LLM during the user’s speaking phase (i.e., the listening phase). The latent reasoning timing label $G$ is similarly applied to ensure that this loss is computed exclusively during the user’s speaking phase.
@@ 5.2 Training pipeline
FLAIR w/o thkthk 0.33 0.49 100 4.3
FLAIR w/ thkthk 0.39 0.46 100 4.3